### v1

# 1st Chenchen Wang College of Information Science

Abstract
This paper develops a reinforcement learning-based controller for the stabilization of partial differential equation (PDE) systems. Within the soft actor-critic (SAC) framework, we embed a DeepONet, a well-known neural operator (NO), which is pretrained using the backstepping controller. The pretrained DeepONet captures the essential features of the backstepping controller and serves as a feature extractor, replacing the convolutional neural networks (CNNs) layers in the original actor and critic networks, and directly connects to the fully connected layers of the SAC architecture. We apply this novel backstepping and reinforcement learning integrated method to stabilize an unstable first-order hyperbolic PDE and an unstable reaction-diffusion PDE. Simulation results demonstrate that the proposed method outperforms the standard SAC, SAC with an untrained DeepONet, and the backstepping controller on both systems.
Index Terms:
Reinforcement Learning, soft actor-critic, DeepONet, first-order hyperbolic PDEs, diffusion-reaction PDEs.
I Introduction
Control of PDE systems remains a highly challenging task due to the infinite-dimensional nature of the state space and the complexity of system dynamics. Learning-based approaches integrated with the classical control theory have emerged as a promising alternative, offering data-driven adaptability [1].
For finite-dimensional systems, there has been a growing body of research [2, 3, 4]. For instance, a control-informed reinforcement learning framework that integrates PID control components into the architecture of deep reinforcement learning (RL) policies is proposed in [5], In this [6], it embeds barrier function theory into neural ordinary differential equations (ODEs), ensuring that the system state remains within a safe region by optimizing the parameters of a learning-based barrier function.
The infinite-dimensional nature of PDE systems poses significant challenges for control design, which has limited research progress in this field. For some special PDE systems there are some RL-based control design, such as applying proximal policy optimization to optimize the boundary controller for the Aw-Rascle-Zhang (ARZ) model of highway traffic flow in [7]. Furthermore, by combining the state-dependent Riccati equation (SDRE) with Bayesian linear regression, [8] proposes a RL-based algorithm for online identification and stable control of unknown PDEs.
However, one of the key challenges in combining classical control theory with learning-based methods lies in effectively incorporating prior knowledge from classical control into neural network learning. To address this issue, we propose a novel approach where a NO, specifically a Deep Operator Network (DeepONet) [9], [10], is pretrained to learn the backstepping controller, a well-established classical control law for PDE systems. DeepONet is particularly suited for this task as it can represent mappings from function spaces to function spaces, making it ideal for approximating the feedback operator. The use of DeepONet to learn backstepping gain functions/controllers is first proposed by [11, 12]. Subsequently, this approach is extended to handle PDE systems with delay [13, 14, 15], applied to enhance the real-time performance of adaptive control of PDEs [16, 17], and utilized in traffic flow control problems [18].
In our method, the DeepONet pretrained on the backstepping controller serves as a feature extractor, replacing the CNN layers in the standard SAC architecture. The DeepONet connects directly to the fully connected layers of the actor and critic networks. During training, the parameters of the DeepONet feature extractor are updated alongside the SAC actor and critic networks, enabling fine-tuning in the RL process.
The primary contribution of this paper is the introduction of a backstepping-pretrained DeepONet into the RL architecture. This accelerates training by providing better initialization, allowing the reward function to start from a higher baseline. We apply this novel backstepping and RL integrated method to stabilize an unstable first-order hyperbolic PDE and an unstable reaction-diffusion PDE. Simulation results show that the proposed method significantly reduces transient oscillations compared to the backstepping controller, vanilla SAC, and SAC with an untrained DeepONet. It also decreases the steady-state error relative to both SAC baseline algorithms. In our experiments, RL-based controllers exhibit small steady-state errors, likely due to the stochastic nature of their policies, whereas the rigorous backstepping controller can eliminate such errors. By integrating a pretrained DeepONet infused with backstepping knowledge, the proposed method markedly reduces this error.
Additionally, when training the DeepONet, we incorporate not only state variables but also system coefficient functions as inputs. This design enables the pretrained DeepONet to adapt its output, the control effort, to variations in system coefficients. As a result, the trained RL controller remains effective even when applied to PDE systems with coefficients different from those seen during training, demonstrating the robustness of the method to parameter variations.
The structure of this paper is as follows: Section II provides problem formulation; Section III presents the NO-based reinforcement learning framework; Section IV validates the effectiveness of the method through simulations; and finally, Section V concludes the paper.
II Problem description
II-A General PDE control problem
We consider a one-dimensional PDE defined on a domain $\mathcal{X}=[0,1]$. The time domain is $\mathcal{T}=[0,T]\subset\mathbb{R}^{+}$. Let $u(x,t)$, where $x\in\mathcal{X}$ and $t\in\mathcal{T}$, describe the state of the system governed by the PDE, with the dynamics given by
$\frac{\partial u}{\partial t}=F\left(u,\frac{\partial u}{\partial x},\frac{\partial^{2}u}{\partial x^{2}},\dots,p_{1}(x),p_{2}(x),\dots U(t)\right),$ (1)
with the initial condition $u(x,0)=u_{0}(x)$, where $F$ is the PDE that models the system dynamics, $U(t)$ is the control function, and $p_{i}(x)$, for $i=1,\dots,m$ are the coefficient functions with $m$ coefficients. The PDE control problem is to design a controller which stabilize the system or a cost function regulates $u(x,t)$ to a desired trajectory.
II-B Backstepping control operator
Employing the PDE backstepping method, one can design a feedback controller, usually the boundary controller, which is generalized by
[…]
where for given $p_{i}(x)\in C[0,1]$ and $u\in C[0,1]$,
[…]
with $k(x,y)$ denoting the backstepping kernel function dependent on the system coefficients.
II-C DeepONet for learning backstepping controller
DeepONet is a deep learning architecture specifically designed for learning “function-to-function” mappings (operators). It consists of two subnetworks: a branch net and a trunk net. The output function is expressed as a weighted sum of basis functions, where the branch net learns the mapping from the input function to the coefficients (which serve as the weights of the basis functions), and the trunk net learns the basis functions themselves.
For the PDE control problem, the inputs to the DeepONet consist of the coefficients $p_{i}(x)$ and state $u(x,\cdot)$, while the output of the DeepONet is the control signal $U(\cdot)$. All training data are generated by solving the PDE (1) with the backstepping controller (2), using different initial conditions and coefficient functions randomly sampled from Chebyshev polynomials with random parameters [11].
The trained DeepONet is adapted for integration into the SAC framework by modifying its output from a control value to a feature vector, as shown in Fig. 1. This feature vector serves as input to the FNN layers of both the actor and critic networks, introducing the backstepping control knowledge into the RL learning.
Figure 1: Architecture of DeepONet as feature extractor for SAC.
III NO-based reinforcement learning framework
In this section, we present the RL framework that integrates a pre-trained DeepONet with the SAC algorithm. We use a SAC framework with an actor network $\pi_{\phi}(a_{t}|s_{t})$, two critic networks $Q_{\theta_{i}}(s_{t},a_{t})$, and their corresponding target networks $Q_{\bar{\theta}_{i}}(s_{t},a_{t})$ (the dual-critic design mitigates bootstrapping issues). The DeepONet introduced in Section II-C, is pre-trained using the backstepping controller, which serves as a feature extractor and replaces the CNN layers in the standard SAC architecture. The DeepONet connects directly to the actor/critic fully connected layers. During training, the actor and critic networks, along with their embedded DeepONet feature extractors, are jointly trained via backpropagation. The architecture of the RL is shown in Fig. 2.
Figure 2: The DeepONet pre-trained with the backstepping method is embedded into the SAC framework.
III-A Actor network
By introducing the policy entropy $H_{\pi}$, it measures the randomness of the action distribution given the state $s_{t}$. The policy entropy is defined as: $H(\pi(\cdot|s_{t}))=-\mathbb{E}_{a_{t}\sim\pi_{\phi}}[\log\pi_{\phi}(a_{t}|s_{t})]$. The policy network $\pi_{\phi}(a_{t}|s_{t})$ is trained by minimizing the following objective function:
$J_{\pi}({\phi})=\mathbb{E}_{s_{t}\sim\mathcal{G}}\mathbb{E}_{a_{t}\sim\pi_{\phi}}\left[-\alpha H(\pi_{\phi}(\cdot|s_{t}))-\min_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})\right],$ (4)
where $\mathcal{G}$ denotes the state distribution. Accordingly, the parameter $\phi$ of the policy network is updated by
$\phi\leftarrow\phi-\eta\nabla_{\phi}J_{\pi}(\phi),$ (5)
where $\eta$ denotes the learning rate.
III-B Critic networks
We use a double Q-network structure (two independent FNNs), the action-value function is defined as $Q_{\pi}(s_{t},a_{t})$, which represents the expected cumulative return that the agent obtains by following the policy $\pi_{\phi}$ after executing the action $a_{t}$ in the state $s_{t}$. Using two target networks $Q_{\bar{\theta}_{1,2}}$ to estimate the target action-value:
$y_{t}=r_{t}+\gamma\left(\min_{i=1,2}Q_{\bar{\theta}_{i}}(s_{t+1},\hat{a}_{t+1})-\alpha\log\pi_{\phi}(\hat{a}_{t+1}|s_{t+1})\right),$ (6)
$\alpha$ is the temperature parameter, which is used to balance the trade-off between reward maximization and policy entropy, and the action $\hat{a}_{t+1}\sim\pi(\cdot|s_{t+1})$ is sampled from the policy network. The weights of the target network $Q_{\bar{\theta}_{i}}$ are updated by
$\bar{\theta}_{i}\leftarrow\tau\theta_{i}+(1-\tau)\bar{\theta}_{i},$ (7)
where $i\in\{1,2\}$ and $\tau\in(0,1]$ denotes the weighted coefficient.
The action-value networks $Q_{\theta_{1}}(s_{t},a_{t})$ and $Q_{\theta_{2}}(s_{t},a_{t})$ are trained by minimizing the following objective:
$J_{Q}(\theta_{i})=\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{W}}\left[(Q_{\theta_{i}}(s_{t},a_{t})-y_{t})^{2}\right],$ (8)
where $i\in\{1,2\}$ and $\mathcal{W}$ is the distribution of the states and actions. The parameter $\theta_{i}$ is updated via stochastic gradient descent:
$\theta_{i}\leftarrow\theta_{i}-\lambda\hat{\nabla}_{\theta_{i}}{J}_{Q}(\theta_{i}),$ (9)
where $\lambda$ denotes the learning rate.
III-C Algorithm (Markov Decision Process)
We have briefly outlined the processing of the Markov Decision Process (MDP), as shown in Algorithm 1.
1) State space $\mathcal{S}$: we denote $s_{t}=\{u(x,t)\}\in\mathcal{S}$ as the augmented state at time $t$.
2) Action space $\mathcal{A}\subseteq\mathbb{R}$: it is defined as $\mathcal{A}=[-\bar{U},\bar{U}]$, where $\bar{U}=\sup_{t\in\mathbb{R}^{+}}|U(t)|$, $U(t)=a_{t}\in\mathcal{A}$.
3) Policy function $\pi(a_{t}|s_{t})$: the policy is a probability density function $\pi(a_{t}|s_{t})$ that maps the current state $s_{t}\in\mathcal{S}$ to a probability distribution over actions $a_{t}\in\mathcal{A}$.
4) State transition $p(s_{t+1}|s_{t},a_{t})$: this $p(s_{t+1}|s_{t},a_{t})$ is the probability distribution of $s_{t+1}$, which depends on the current state $s_{t}$ and the action $a_{t}$ chosen by the agent. Given the state $s_{t}$ at time $t$, when the agent executes action $a_{t}$, the environment transitions to the next state $s_{t+1}$ following this probability distribution.
5) Reward function $r_{t}$ and additional reward function $q$: it is designed to guide the learning of the optimal control policy. The reward function is defined as follows:
$r_{t}(s_{t},s_{t+1})$ $=-1\cdot\|s_{t+1}-s_{t}\|_{L_{2}},$ (10)
$q(s_{T},a_{0},\ldots,a_{T})$ =$=\begin{cases}0&\|s_{T}\|_{L_{2}}>\zeta,\\ \begin{aligned} \sigma-\dfrac{1}{\eta}\cdot\sum_{\tau=0}^{T}|a_{\tau}|_{L_{1}}\\ -\|s_{t}\|_{L_{2}}\end{aligned}&\|s_{T}\|_{L_{2}}\leq\zeta,\end{cases}$ (11)
here, $\sigma$, $\eta$, and $\zeta$ are hyperparameters.
The reward function $r_{t}$ is applied at each step to drive state convergence, while an additional reward $q$ is given at the end of each episode if the $L_{2}$ norm of the final state is less than or equal to a specified threshold $\zeta$.
Algorithm 1 SAC Procedure
1:Initialization: $\phi,\theta_{i},\bar{\theta}_{i}$ for $i=1,2$, $\mathcal{W}$ and $s_{0}=\{u_{0}(x)\}$
2:for $l\leftarrow 0$ to total steps do
3:     if $\lVert s_{t}\rVert_{L_{2}}>\mathrm{limit}\lor t\geq T$ then
4:         Reset initial state $s_{0}$
5:     end if
6:     Observe current state $s_{t}=\{u_{t}\}$
7:     Sample action $a_{t}\sim\pi_{\phi}(a_{t}\mid s_{t})$
8:     Execute $a_{t}$, observe $s_{t+1}$ and $r_{t}$
9:     $\mathcal{W}\leftarrow\mathcal{W}\cup\{(s_{t},a_{t},r_{t},s_{t+1})\}$
10:     for each gradient step do
11:         Sample a mini-batch from $\mathcal{W}$
12:         Compute the target value $y_{t}$ using (6)
13:         Update parameters of the $Q_{\theta_{1,2}}$ by (8)
14:         Update parameters of the $Q_{\bar{\theta}_{1,2}}$ by (5)
15:         Update parameters of the $\pi_{\phi}$ by (4)
16:     end for
17:end for
IV Simulation
In this section, we apply the proposed RL controller to both a hyperbolic PDE and a parabolic PDE. Details of the corresponding backstepping designs can be found in [19]. The standard SAC code is obtained from the PDE Control Gym platform introduced in [20]. We compare our method (NOSAC_training) with the backstepping controller (Backstepping), the standard SAC (SAC), and SAC embedded with DeepONet without pretraining the DeepONet using the backstepping controller (NOSAC).
Note that all the experiments are carried out on a workstation equipped with an Intel Core i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU. For all the hyperparameters required by the SAC algorithm and the hyperparameters of the reward function, please refer to Appendix A.2 of [20].
IV-A Simulation of a 1D Hyperbolic PDE
First, we train a DeepONet to approximate the backstepping controller. To generate the training data. We employ the finite difference method with a spatial step size of $dx=0.01$ over the time interval $[0,5]$ with a temporal step size of $\Delta t=0.001$. The coefficient $\beta(x)$ is sampled from the Chebyshev polynomial $\beta(x)=5\cos\big{(}\gamma\cos^{-1}(x)\big{)}$, where $\gamma\sim U[5.5,7]$. The initial condition $u_{0}(x)$ is assumed to be a constant randomly chosen from the uniform distribution $U[1,10]$.
We generate a dataset comprising $6\times 10^{5}$ samples, which were produced by combining 100 different $\beta(x)$ and 60 different $u_{0}(x)$. During the data collection process, data was collected every 50 time steps. After completing the entire data collection, we randomly shuffled the dataset and split it into a training set 90% and a testing set 10%.
Figure 3: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for hyperbolic PDE with $\gamma=5.5$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
The DeepONet approximation results are shown in Fig. 3, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=5.5$. It can be seen that the state converges to zero under both controllers.
Figure 4: Reward curves for the three RL training processes: (a) 1D Hyperbolic PDE, (b) 1D Parabolic PDE.
The same $\gamma$ value is used, and the initial state $u_{0}(x)$ is randomly selected during RL. Training consists of 100,000 steps, with 100 interaction steps per episode. Fig. 4(a) shows the reward curves for SAC, NOSAC, and NOSAC_training on 1D hyperbolic PDE, where NOSAC_training achieves faster reward growth and finds effective policies more quickly than the other two.
We apply the NOSAC_training controller to the 1D hyperbolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig. 5. The backstepping controller yields smoother closed-loop state evolution, while the proposed NOSAC_training controller exhibits less overshoot. To better illustrate performance, Fig. 6 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training outperforms the backstepping controller, SAC, and NOSAC in terms of overshoot and convergence rate, while the backstepping controller achieves lower steady-state error.
To evaluate robustness under model mismatch, we simulate a system with $\gamma=5.7$ using the three LR_based controllers trained at $\gamma=5.5$ and the backstepping controller designed for $\gamma=5.5$. The results in Fig. 6 demonstrate that NOSAC_training achieves superior robustness, outperforming the other controllers in overshoot, convergence rate, and steady-state error.
Figure 5: Closed-loop state evolution with $\gamma=5.5$ and $u_{0}(x)=9$ under (a) Backstepping, (b) SAC, (c) NOSAC, (d) NOSAC_training.
Figure 6: (a) Control input and (b) $L_{2}$ norm of the state with $\gamma=5.5$ and $u_{0}(x)=9$ under the four different controllers.
Figure 7: (a) Control input and (b) $L_{2}$ norm of the state under model mismatch with system’s parameter $\gamma=5.7$ and initial condition $u_{0}(x)=9$ while training the three RLs and designing the backstepping using $\gamma=5.5$.
IV-B Simulation of a 1D Parabolic PDE
When dealing with 1D Parabolic PDE, the implicit method is employed, which enhances accuracy and reduces computation. We employ the $dx=0.01$ over the time interval $[0,1]$ with a temporal step size of $\Delta t=0.001$. The coefficient $\lambda(x)$ is sampled from the Chebyshev polynomial $\lambda(x)=50\cos\big{(}\gamma\cos^{-1}(x)\big{)}$, where $\gamma\sim U[8,12]$. We generate a dataset comprising $6\times 10^{5}$ samples, which were produced by combining 100 different $\lambda(x)$ and 60 different $u_{0}(x)$. During the data collection process, data was collected every 100 time steps, its DeepONet training process is the same as that for the 1D Hyperbolic PDE. During testing, we extend the time $T$ to 3 seconds to observe the effects.
The DeepONet approximation results are shown in Fig. 8, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=9$. It can be seen that the state converges to zero under both controllers.
Figure 8: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for hyperbolic PDE with $\gamma=9$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
The same $\gamma$ value is used, and the initial state $u_{0}(x)$ is randomly selected during the RL, the RL training process is the same as that for 1D Hyperbolic PDE. Fig. 4(b) shows the reward curves for SAC, NOSAC, and NOSAC_training on 1D Parabolic PDE, where NOSAC_training achieves faster reward growth.
We apply the NOSAC_training controller to the 1D parabolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig. 9. The proposed NOSAC_training controller exhibits less overshoot. Fig. 10 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training is optimal in terms of overshoot and convergence speed, while the backstepping controller achieves a lower steady-state error.
Figure 9: Closed-loop state evolution with $\gamma=9$ and $u_{0}(x)=9$ under (a) Backstepping, (b) SAC, (c) NOSAC, (d) NOSAC_training.
Figure 10: (a) Control input and (b) $L_{2}$ norm of the state with $\gamma=9$ and $u_{0}(x)=9$ under the four different controllers.
To evaluate robustness under model mismatch, we simulate a system with $\gamma=8.5$ using the three LR_based controllers trained at $\gamma=9$ and the backstepping controller designed for $\gamma=9$. The results in Fig. 11 demonstrate that NOSAC_training achieves superior robustness, outperforming the other controllers in overshoot, convergence rate, and steady-state error. Note that the backtepping controller performs worse for the parabolic PDE than the hyperbolic PDE under model mismatch.
Figure 11: (a) Control input and (b) $L_{2}$ norm of the state under model mismatch with system’s parameter $\gamma=8.5$ and initial condition $u_{0}(x)=9$ while training the three RLs and designing the backstepping using $\gamma=9$.
V Conclusion
This paper develops a novel neural-operator-embedded SAC architecture for PDE control, which incorporates the knowledge of rigorous control through a DeepONet that approximates the backstepping controller. We apply this learning-based control methodology to both a 1D hyperbolic and a 1D parabolic unstable PDE, and evaluate its performance in terms of overshoot, convergence rate, and steady-state error, compared to the backstepping controller, standard SAC, and SAC embedded with a non-pretrained DeepONet. Simulations on both systems demonstrate that the proposed method outperforms the other three controllers. The robustness against model mismatch is also evaluated by changing the system coefficient to a value different from that used for control design. The results show that the proposed method achieves superior performance compared to the other approaches, owing to the pre-trained DeepONet’s ability to adapt to varying coefficient functions and generate appropriate control signals in response to these changes. Future research will explore integrating safety control with this learning-based approach.

### v2

@@
-1st Chenchen Wang College of Information Science
+1st Chenchen Wang    2nd Jie Qi    3rd Jiaqi Hu
 Abstract
@@
 Within the Soft Actor-Critic (SAC) framework, we embed a DeepONet, a well-known neural operator (NO), which is pretrained using the backstepping controller.
-The pretrained DeepONet captures the essential features of the backstepping controller and serves as a feature extractor, replacing the convolutional neural networks (CNNs) layers in the original actor and critic networks, and directly connects to the fully connected layers of the SAC architecture.
-We apply this novel backstepping and reinforcement learning integrated method to stabilize an unstable first-order hyperbolic PDE and an unstable reaction-diffusion PDE.
-Simulation results demonstrate that the proposed method outperforms the standard SAC, SAC with an untrained DeepONet, and the backstepping controller on both systems.
-Index Terms:
-Reinforcement Learning, soft actor-critic, DeepONet, first-order hyperbolic PDEs, diffusion-reaction PDEs.
+The pretrained DeepONet captures the essential features of the backstepping controller and is used to extract features, replacing the convolutional neural networks (CNNs) in the original actor-critic networks, and directly connects to the fully connected layers of the SAC architecture.
+We apply this novel backstepping and reinforcement learning integrated method to stabilize an unstable 1D hyperbolic PDE and an unstable reaction-diffusion PDE.
+The proposed method is shown to outperform standard SAC in simulations, SAC with an untrained DeepONet, and the backstepping controller on both systems.
 I Introduction
@@
 For finite-dimensional systems, there has been a growing body of research [2, 3, 4].
-For instance, a control-informed reinforcement learning framework that integrates PID control components into the architecture of deep reinforcement learning (RL) policies is proposed in [5], In this [6], it embeds barrier function theory into neural ordinary differential equations (ODEs), ensuring that the system state remains within a safe region by optimizing the parameters of a learning-based barrier function.
+For instance, a control-informed reinforcement learning (RL) framework that integrates PID control components into the architecture of deep RL policies is proposed in [5], In this [6], it embeds barrier function theory into neural ordinary differential equations (ODEs), ensuring that the system state remains within a safe region by optimizing the parameters of a learning-based barrier function.
 The infinite-dimensional nature of PDE systems poses significant challenges for control design, which has limited research progress in this field.
@@
 Subsequently, this approach is extended to handle PDE systems with delay [13, 14, 15], applied to enhance the real-time performance of adaptive control of PDEs [16, 17], and utilized in traffic flow control problems [18].
-In our method, the DeepONet pretrained on the backstepping controller serves as a feature extractor, replacing the CNN layers in the standard SAC architecture.
-The DeepONet connects directly to the fully connected layers of the actor and critic networks.
-During training, the parameters of the DeepONet feature extractor are updated alongside the SAC actor and critic networks, enabling fine-tuning in the RL process.
+In our method, the DeepONet pretrained on the backstepping controller is used for feature extraction, as a replacement for the CNNs in standard SAC.
+The DeepONet connects directly to the fully connected actor-critic networks layers.
+Throughout the training phase, parameters of the DeepONet feature extraction module are optimized in conjunction with those of the SAC actor-critic networks, thereby enabling fine-tuning during reinforcement learning.
 The primary contribution of this paper is the introduction of a backstepping-pretrained DeepONet into the RL architecture.
@@
 As a result, the trained RL controller remains effective even when applied to PDE systems with coefficients different from those seen during training, demonstrating the robustness of the method to parameter variations.
-The structure of this paper is as follows: Section II provides problem formulation; Section III presents the NO-based reinforcement learning framework; Section IV validates the effectiveness of the method through simulations; and finally, Section V concludes the paper.
+The structure of this paper is as follows: Section II provides problem formulation; Section III presents the NO-based RL framework; Section IV validates the effectiveness of the method through simulations; and finally, Section V concludes the paper.
 II Problem description
@@
 $\frac{\partial u}{\partial t}=F\left(u,\frac{\partial u}{\partial x},\frac{\partial^{2}u}{\partial x^{2}},\dots,p_{1}(x),p_{2}(x),\dots U(t)\right),$ (1)
-with the initial condition $u(x,0)=u_{0}(x)$, where $F$ is the PDE that models the system dynamics, $U(t)$ is the control function, and $p_{i}(x)$, for $i=1,\dots,m$ are the coefficient functions with $m$ coefficients.
-The PDE control problem is to design a controller which stabilize the system or a cost function regulates $u(x,t)$ to a desired trajectory.
+with the initial condition $u(x,0)=u_{0}(x)$, where $F$ is the PDE that models the system dynamics, $U(t)$ is the control function, and $p_{i}(x)$, for $i=1,\dots,m$ are the coefficient functions.
+The PDE control problem is to design a controller which stabilizes the system or regulates states to a desired trajectory.
 II-B Backstepping control operator
-Employing the PDE backstepping method, one can design a feedback controller, usually the boundary controller, which is generalized by
-[…]
-where for given $p_{i}(x)\in C[0,1]$ and $u\in C[0,1]$,
-[…]
-with $k(x,y)$ denoting the backstepping kernel function dependent on the system coefficients.
+Employing the PDE backstepping method [19], a feedback controller—typically implemented as a boundary controller—can be designed in the following form:
+$U(t)$ $=\int_{0}^{1}k(1,y)u(y,t)\,dy,$ (2)
+where $k(1,y)$ is the control gain derived from the backstepping kernel function.
+This kernel depends on the system coefficient functions $p_{i}(x)$, $i=1,\ldots,m$.
+To formalize the controller structure, we define the control operator $\mathcal{U}:(C[0,1])^{m}\times L^{2}[0,1]\mapsto\mathbb{R}$, which maps the coefficient functions $(p_{1},\ldots,p_{m})\in(C[0,1])^{m}$ and the system state $u\in L^{2}[0,1]$ to the control input $U$.
+Specifically,
+$U:=\mathcal{U}(p_{1},\dots,p_{m},u),$ (3)
+where $U$ is defined as in (2).
 II-C DeepONet for learning backstepping controller
-DeepONet is a deep learning architecture specifically designed for learning “function-to-function” mappings (operators).
-It consists of two subnetworks: a branch net and a trunk net.
-The output function is expressed as a weighted sum of basis functions, where the branch net learns the mapping from the input function to the coefficients (which serve as the weights of the basis functions), and the trunk net learns the basis functions themselves.
-For the PDE control problem, the inputs to the DeepONet consist of the coefficients $p_{i}(x)$ and state $u(x,\cdot)$, while the output of the DeepONet is the control signal $U(\cdot)$.
+DeepONet is a deep learning architecture designed to learn function-to-function mappings, consisting of branch and trunk networks.
+The output function is a weighted sum of basis functions, with the branch network learning the weights and the trunk network learning the basis functions.
+For the PDE controller approximated by a DeepONet, the DeepONet receives coefficient functions $p_{i}(x)$ and state $u(x,\cdot)$ as inputs, generating a scalar control signal $U(\cdot)$ as output.
+In the mapping, the input and output are time-independent, the control input is generated based on the current system state.
 All training data are generated by solving the PDE (1) with the backstepping controller (2), using different initial conditions and coefficient functions randomly sampled from Chebyshev polynomials with random parameters [11].
-The trained DeepONet is adapted for integration into the SAC framework by modifying its output from a control value to a feature vector, as shown in Fig.
+To train the DeepONet, we employ the following state-based loss function:
+$\varepsilon=|U_{NO}-U|,$ (4)
+where $U_{NO}$ denotes the control generated by DeepONet, while $U$ denotes the backstepping control.
+To integrate the DeepONet into the reinforcement learning architecture, we modify its output from a scalar control signal to a vector representation, enabling feature extraction for both actor-critic networks in SAC, as depicted in Fig.
 1.
-This feature vector serves as input to the FNN layers of both the actor and critic networks, introducing the backstepping control knowledge into the RL learning.
+The dimensionality of the DeepONet output is determined by its branch and trunk networks, along with subsequent fully connected layers.
+The pre-trained DeepONet encapsulates the knowledge of the backstepping controller, enabling seamless integration of prior control strategies into the RL decision-making process.
+This incorporation not only improves the training efficiency of the reinforcement learning algorithm but also enhances the performance of the RL-based controller.
 Figure 1: Architecture of DeepONet as feature extractor for SAC.
 III NO-based reinforcement learning framework
-In this section, we present the RL framework that integrates a pre-trained DeepONet with the SAC algorithm.
-We use a SAC framework with an actor network $\pi_{\phi}(a_{t}|s_{t})$, two critic networks $Q_{\theta_{i}}(s_{t},a_{t})$, and their corresponding target networks $Q_{\bar{\theta}_{i}}(s_{t},a_{t})$ (the dual-critic design mitigates bootstrapping issues).
-The DeepONet introduced in Section II-C, is pre-trained using the backstepping controller, which serves as a feature extractor and replaces the CNN layers in the standard SAC architecture.
-The DeepONet connects directly to the actor/critic fully connected layers.
-During training, the actor and critic networks, along with their embedded DeepONet feature extractors, are jointly trained via backpropagation.
-The architecture of the RL is shown in Fig.
+In this section, we elaborate on the reinforcement learning (RL) framework that fuses a pre-trained DeepONet with the Soft Actor-Critic (SAC) algorithm.
+This framework is composed of an actor network $\pi_{\phi}(a_{t}|s_{t})$, a pair of critic networks denoted as $Q_{\theta_{i}}(s_{t},a_{t})$, as well as their respective target networks $Q_{\bar{\theta}_{i}}(s_{t},a_{t})$.
+Serving as a feature extraction component, the DeepONet replaces the CNNs adopted in the canonical SAC structure, and establishes a direct connection with the fully connected layers of the actor and critic networks.
+In the training phase, the actor-critic networks and the embedded DeepONet feature extractor are optimized in a joint manner through the backpropagation algorithm.
+The schematic diagram of this RL framework is presented in Fig.
 2.
-Figure 2: The DeepONet pre-trained with the backstepping method is embedded into the SAC framework.
 III-A Actor network
-By introducing the policy entropy $H_{\pi}$, it measures the randomness of the action distribution given the state $s_{t}$.
+By introducing the policy entropy $H_{\pi}$, it quantifies the uncertainty of the action distribution conditioned on state $s_{t}$.
 The policy entropy is defined as: $H(\pi(\cdot|s_{t}))=-\mathbb{E}_{a_{t}\sim\pi_{\phi}}[\log\pi_{\phi}(a_{t}|s_{t})]$.
-The policy network $\pi_{\phi}(a_{t}|s_{t})$ is trained by minimizing the following objective function:
-$J_{\pi}({\phi})=\mathbb{E}_{s_{t}\sim\mathcal{G}}\mathbb{E}_{a_{t}\sim\pi_{\phi}}\left[-\alpha H(\pi_{\phi}(\cdot|s_{t}))-\min_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})\right],$ (4)
-where $\mathcal{G}$ denotes the state distribution. Accordingly, the parameter $\phi$ of the policy network is updated by
-$\phi\leftarrow\phi-\eta\nabla_{\phi}J_{\pi}(\phi),$ (5)
-where $\eta$ denotes the learning rate.
+The policy network $\pi_{\phi}(a_{t}|s_{t})$ is trained by optimizing the following subsequent objective function:
+$J_{\pi}({\phi})=\mathbb{E}_{s_{t}\sim\mathcal{D}}\mathbb{E}_{a_{t}\sim\pi_{\phi}}\left[-\alpha H(\pi_{\phi}(\cdot|s_{t}))-\min_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})\right],$ (5)
+with $\mathcal{D}$ characterizes the distribution of system states. Consequently, the update of the policy network parameter $\phi$ is realized by
+$\phi\leftarrow\phi-\eta\nabla_{\phi}J_{\pi}(\phi).$ (6)
 III-B Critic networks
 We use a double Q-network structure (two independent FNNs), the action-value function is defined as $Q_{\pi}(s_{t},a_{t})$, which represents the expected cumulative return that the agent obtains by following the policy $\pi_{\phi}$ after executing the action $a_{t}$ in the state $s_{t}$.
-Using two target networks $Q_{\bar{\theta}_{1,2}}$ to estimate the target action-value:
+Using two target networks $Q_{\bar{\theta}_{1,2}}$ to determine the target Q-function:
 $y_{t}=r_{t}+\gamma\left(\min_{i=1,2}Q_{\bar{\theta}_{i}}(s_{t+1},\hat{a}_{t+1})-\alpha\log\pi_{\phi}(\hat{a}_{t+1}|s_{t+1})\right),$ (7)
-$\alpha$ is the temperature parameter, which is used to balance the trade-off between reward maximization and policy entropy, and the action $\hat{a}_{t+1}\sim\pi(\cdot|s_{t+1})$ is sampled from the policy network.
+$\alpha$ is the temperature parameter for balancing reward maximization and policy entropy, and the action $\hat{a}_{t+1}\sim\pi(\cdot|s_{t+1})$ is drawn from the actor network.
 The weights of the target network $Q_{\bar{\theta}_{i}}$ are updated by
@@
 $J_{Q}(\theta_{i})=\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{W}}\left[(Q_{\theta_{i}}(s_{t},a_{t})-y_{t})^{2}\right],$ (9)
-where $i\in\{1,2\}$ and $\mathcal{W}$ is the distribution of the states and actions. The parameter $\theta_{i}$ is updated via stochastic gradient descent:
-$\theta_{i}\leftarrow\theta_{i}-\lambda\hat{\nabla}_{\theta_{i}}{J}_{Q}(\theta_{i}),$ (9)
-where $\lambda$ denotes the learning rate.
+where $i\in\{1,2\}$ and $\mathcal{W}$ is the distribution of the states and actions. The parameter $\theta_{i}$ is updated by stochastic gradient descent:
+$\theta_{i}\leftarrow\theta_{i}-\lambda\hat{\nabla}_{\theta_{i}}{J}_{Q}(\theta_{i}).$ (10)
+Figure 2: The DeepONet pre-trained with the backstepping method is embedded into the SAC framework.
 III-C Algorithm (Markov Decision Process)
 We have briefly outlined the processing of the Markov Decision Process (MDP), as shown in Algorithm 1.
-1) State space $\mathcal{S}$: we denote $s_{t}=\{u(x,t)\}\in\mathcal{S}$ as the augmented state at time $t$.
-2) Action space $\mathcal{A}\subseteq\mathbb{R}$: it is defined as $\mathcal{A}=[-\bar{U},\bar{U}]$, where $\bar{U}=\sup_{t\in\mathbb{R}^{+}}|U(t)|$, $U(t)=a_{t}\in\mathcal{A}$.
-3) Policy function $\pi(a_{t}|s_{t})$: the policy is a probability density function $\pi(a_{t}|s_{t})$ that maps the current state $s_{t}\in\mathcal{S}$ to a probability distribution over actions $a_{t}\in\mathcal{A}$.
+1) State space $\mathcal{S}$: we denote $s_{t}=\{u(x,t)\}\in\mathcal{S}$, which constitutes the extended state at time $t$.
+2) Action space $\mathcal{A}\subseteq\mathbb{R}$: it is $\mathcal{A}=[-\bar{U},\bar{U}]$, with $\bar{U}=\sup_{t\in\mathbb{R}^{+}}|U(t)|$, $U(t)=a_{t}\in\mathcal{A}$.
+3) Policy function $\pi(a_{t}|s_{t})$: this policy is formulated as a probability density function $\pi(a_{t}|s_{t})$, which establishes a mapping relationship from the current state space $\mathcal{S}$ to the probability distribution over the action space $\mathcal{A}$.
 4) State transition $p(s_{t+1}|s_{t},a_{t})$: this $p(s_{t+1}|s_{t},a_{t})$ is the probability distribution of $s_{t+1}$, which depends on the current state $s_{t}$ and the action $a_{t}$ chosen by the agent.
-Given the state $s_{t}$ at time $t$, when the agent executes action $a_{t}$, the environment transitions to the next state $s_{t+1}$ following this probability distribution.
 5) Reward function $r_{t}$ and additional reward function $q$: it is designed to guide the learning of the optimal control policy. The reward function is defined as follows:
 $r_{t}(s_{t},s_{t+1})$ $=-1\cdot\|s_{t+1}-s_{t}\|_{L_{2}},$ (11)
-$q(s_{T},a_{0},\ldots,a_{T})$ =$=\begin{cases}0&\|s_{T}\|_{L_{2}}>\zeta,\\ \begin{aligned} \sigma-\dfrac{1}{\eta}\cdot\sum_{\tau=0}^{T}|a_{\tau}|_{L_{1}}\\ -\|s_{t}\|_{L_{2}}\end{aligned}&\|s_{T}\|_{L_{2}}\leq\zeta,\end{cases}$ (11)
+$q(s_{T},a_{0},\ldots,a_{T})$ $=\begin{cases}0&\|s_{T}\|_{L_{2}}>\zeta,\\ \begin{aligned} \sigma-\dfrac{1}{\eta}\cdot\sum_{\tau=0}^{T}|a_{\tau}|_{L_{1}}\\ -\|s_{t}\|_{L_{2}}\end{aligned}&\|s_{T}\|_{L_{2}}\leq\zeta,\end{cases}$ (12)
 here, $\sigma$, $\eta$, and $\zeta$ are hyperparameters.
-The reward function $r_{t}$ is applied at each step to drive state convergence, while an additional reward $q$ is given at the end of each episode if the $L_{2}$ norm of the final state is less than or equal to a specified threshold $\zeta$.
-Algorithm 1 SAC Procedure
-1:Initialization: $\phi,\theta_{i},\bar{\theta}_{i}$ for $i=1,2$, $\mathcal{W}$ and $s_{0}=\{u_{0}(x)\}$
-2:for $l\leftarrow 0$ to total steps do
-3:     if $\lVert s_{t}\rVert_{L_{2}}>\mathrm{limit}\lor t\geq T$ then
-4:         Reset initial state $s_{0}$
+The reward function $r_{t}$ drives state convergence at each step, with an additional reward $q$ given at the end of each episode if the $L_{2}$ norm of the final state is less than or equal to the threshold $\zeta$.
+Algorithm 1 DeepONet-SAC Training Process
+1:Initialize network parameters: $\phi,\theta_{i},\bar{\theta}_{i}$ ($i=1,2$), replay memory $\mathcal{W}$, and starting state $s_{0}=\{u_{0}(x)\}$
+2:for iteration $l$ from 0 to max steps do
+3:  if $\|s_{t}\|_{L_{2}}>\text{threshold}$ or $t\geq T$ then
+4:   Reinitialize state $s_{t}\leftarrow s_{0}$
 5:  end if
-6:     Observe current state $s_{t}=\{u_{t}\}$
-7:     Sample action $a_{t}\sim\pi_{\phi}(a_{t}\mid s_{t})$
-8:     Execute $a_{t}$, observe $s_{t+1}$ and $r_{t}$
-9:     $\mathcal{W}\leftarrow\mathcal{W}\cup\{(s_{t},a_{t},r_{t},s_{t+1})\}$
-10:     for each gradient step do
-11:         Sample a mini-batch from $\mathcal{W}$
-12:         Compute the target value $y_{t}$ using (6)
-13:         Update parameters of the $Q_{\theta_{1,2}}$ by (8)
-14:         Update parameters of the $Q_{\bar{\theta}_{1,2}}$ by (5)
-15:         Update parameters of the $\pi_{\phi}$ by (4)
+6:  Set the instantaneous state $s_{t}=\{u_{t}\}$
+7:  Draw action $a_{t}\sim\pi_{\phi}(\cdot\mid s_{t})$
+8:  Apply $a_{t}$, observe next state $s_{t+1}$ and reward $r_{t}$
+9:  Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{W}$
+10:  for each gradient update iteration do
+11:   Randomly sample mini-batch from $\mathcal{W}$
+12:   Calculate target value $y_{t}$ via equation (6)
+13:   Adjust critic parameters $Q_{\theta_{1,2}}$ using (8)
+14:   Update target network parameters $Q_{\bar{\theta}_{1,2}}$ by (5)
+15:   Modify actor parameters $\phi$ according to (4)
 16:  end for
@@
 IV Simulation
-In this section, we apply the proposed RL controller to both a hyperbolic PDE and a parabolic PDE.
-Details of the corresponding backstepping designs can be found in [19].
-The standard SAC code is obtained from the PDE Control Gym platform introduced in [20].
-We compare our method (NOSAC_training) with the backstepping controller (Backstepping), the standard SAC (SAC), and SAC embedded with DeepONet without pretraining the DeepONet using the backstepping controller (NOSAC).
-Note that all the experiments are carried out on a workstation equipped with an Intel Core i9-13900KF CPU and an NVIDIA GeForce RTX 4090 GPU.
-For all the hyperparameters required by the SAC algorithm and the hyperparameters of the reward function, please refer to Appendix A.2 of [20].
+In this section, the proposed RL controller is applied to both hyperbolic and parabolic PDEs, and the details of the backstepping design can be found in [19].
+We adopt the standard SAC implementation from the PDE Control Gym platform introduced in [20].
+Our proposed method—SAC embedded with DeepONet pre-trained using the backstepping controller (referred to as NOSAC_training)—is compared against four baselines: the backstepping controller (Backstepping), standard SAC (SAC), and SAC embedded with DeepONet without pretraining (NOSAC).
+Experiments conducted on an Intel i9-13900KF/RTX 4090 workstation; SAC and reward hyperparameters in [20] Appendix A.2.
 IV-A Simulation of a 1D Hyperbolic PDE
-First, we train a DeepONet to approximate the backstepping controller.
-To generate the training data.
-We employ the finite difference method with a spatial step size of $dx=0.01$ over the time interval $[0,5]$ with a temporal step size of $\Delta t=0.001$.
+First, A DeepONet is trained to emulate the backstepping controller, with corresponding training data generated accordingly.
+We employ the finite difference method with $dx=0.01$ over the time interval $[0,5]$ with a temporal step size of $\Delta t=0.001$.
 The coefficient $\beta(x)$ is sampled from the Chebyshev polynomial $\beta(x)=5\cos\big(\gamma\cos^{-1}(x)\big)$, where $\gamma\sim U[5.5,7]$.
 The initial condition $u_{0}(x)$ is assumed to be a constant randomly chosen from the uniform distribution $U[1,10]$.
-We generate a dataset comprising $6\times 10^{5}$ samples, which were produced by combining 100 different $\beta(x)$ and 60 different $u_{0}(x)$.
-During the data collection process, data was collected every 50 time steps.
-After completing the entire data collection, we randomly shuffled the dataset and split it into a training set 90% and a testing set 10%.
-Figure 3: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for hyperbolic PDE with $\gamma=5.5$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
+We generate a dataset of $6\times 10^{5}$ samples, which comprises 100 different $\beta(x)$ and 60 different $u_{0}(x)$.
+We collect data every 50 time steps, and then randomly shuffle and split the dataset into a training set 90% and a testing set 10%.
+The training process requires approximately 175 minutes.
 The DeepONet approximation results are shown in Fig.
 3, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=5.5$.
-It can be seen that the state converges to zero under both controllers.
+It can be seen that $U_{NO}$ remains almost indistinguishable from $U$, and the state converges to zero under both controllers ($\|u_{NO}-u\|_{L_{2}}$ is less than $10^{-4}$ after about $3$ seconds).
+Figure 3: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for hyperbolic PDE with $\gamma=5.5$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
 Figure 4: Reward curves for the three RL training processes: (a) 1D Hyperbolic PDE, (b) 1D Parabolic PDE.
-The same $\gamma$ value is used, and the initial state $u_{0}(x)$ is randomly selected during RL.
-Training consists of 100,000 steps, with 100 interaction steps per episode.
+In the RL training, we set $\gamma=5.5$ and sample $u_{0}(x)$ sampled from a uniform distribution $U[1,10]$ at each iteration.
+The training consists of 100,000 steps with 100 interaction steps per episode.
+The training time is about 18, 20, and 11 minutes for NOSAC_training, NOSAC, and SAC taking, respectively.
 Fig.
-4(a) shows the reward curves for SAC, NOSAC, and NOSAC_training on 1D hyperbolic PDE, where NOSAC_training achieves faster reward growth and finds effective policies more quickly than the other two.
+4(a) shows NOSAC_training achieves faster reward growth and policy convergence compared to the others, thanks to the pre-trained DeepONet.
+Pre-trained on backstepping controller, the DeepONet captures stable control behaviors, enabling it to encode this knowledge into feature vectors with higher fidelity.
+This serves as a warm start, for the SAC actor—reducing exploration costs and increasing initial rewards—and provides the critic with more informative state features.
 We apply the NOSAC_training controller to the 1D hyperbolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig.
@@
 To better illustrate performance, Fig.
-6 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training outperforms the backstepping controller, SAC, and NOSAC in terms of overshoot and convergence rate, while the backstepping controller achieves lower steady-state error.
+6 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training outperforms the other controllers in terms of overshoot and convergence rate, whereas the backstepping controller demonstrates superior steady-state precision.
 To evaluate robustness under model mismatch, we simulate a system with $\gamma=5.7$ using the three LR_based controllers trained at $\gamma=5.5$ and the backstepping controller designed for $\gamma=5.5$.
 The results in Fig.
-6 demonstrate that NOSAC_training achieves superior robustness, outperforming the other controllers in overshoot, convergence rate, and steady-state error.
+7 demonstrate that NOSAC_training achieves superior robustness, outperforming the SAC and NOSAC controllers in overshoot, convergence speed, and steady-state error.
 Figure 5: Closed-loop state evolution with $\gamma=5.5$ and $u_{0}(x)=9$ under (a) Backstepping, (b) SAC, (c) NOSAC, (d) NOSAC_training.
@@
 During the data collection process, data was collected every 100 time steps, its DeepONet training process is the same as that for the 1D Hyperbolic PDE.
-During testing, we extend the time $T$ to 3 seconds to observe the effects.
+The total training process takes about 190 minutes.
 The DeepONet approximation results are shown in Fig.
-8, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=9$.
-It can be seen that the state converges to zero under both controllers.
-Figure 8: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for hyperbolic PDE with $\gamma=9$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
-The same $\gamma$ value is used, and the initial state $u_{0}(x)$ is randomly selected during the RL, the RL training process is the same as that for 1D Hyperbolic PDE.
+8, illustrating the boundary control and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=9$.
+It can be seen that $U_{NO}$ remains almost indistinguishable from $U$, and the state converges to zero under both controllers($\|u_{NO}-u\|_{L_{2}}$ is less than $10^{-4}$ after about $0.5$ seconds).
+Figure 8: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for parabolic PDE with $\gamma=9$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
+In the RL training, we set $\gamma=9$, while the other settings are consistent with those used in the hyperbolic PDE.
+The training processes of NOSAC_training, NOSAC, and SAC take approximately 20 minutes, 22 minutes, and 14 minutes, respectively.
 Fig.
-4(b) shows the reward curves for SAC, NOSAC, and NOSAC_training on 1D Parabolic PDE, where NOSAC_training achieves faster reward growth.
+4(b) shows the reward curves for the proposed method and other two baseline algorithms on 1D Parabolic PDE, demonstrating that NOSAC_training exhibits the fastest reward growth and convergence.
 We apply the NOSAC_training controller to the 1D parabolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig.
@@
 Fig.
-10 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training is optimal in terms of overshoot and convergence speed, while the backstepping controller achieves a lower steady-state error.
+10 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training is optimal in terms of overshoot and convergence speed, conversely, the backstepping controller exhibits enhanced steady-state convergence.
 Figure 9: Closed-loop state evolution with $\gamma=9$ and $u_{0}(x)=9$ under (a) Backstepping, (b) SAC, (c) NOSAC, (d) NOSAC_training.
@@
 The results in Fig.
-11 demonstrate that NOSAC_training achieves superior robustness, outperforming the other controllers in overshoot, convergence rate, and steady-state error.
-Note that the backtepping controller performs worse for the parabolic PDE than the hyperbolic PDE under model mismatch.
+11 demonstrate that NOSAC_training achieves superior robustness, outperforming the SAC and NOSAC controllers in overshoot, convergence speed, and steady-state error.
 Figure 11: (a) Control input and (b) $L_{2}$ norm of the state under model mismatch with system’s parameter $\gamma=8.5$ and initial condition $u_{0}(x)=9$ while training the three RLs and designing the backstepping using $\gamma=9$.
 V Conclusion
-This paper develops a novel neural-operator-embedded SAC architecture for PDE control, which incorporates the knowledge of rigorous control through a DeepONet that approximates the backstepping controller.
-We apply this learning-based control methodology to both a 1D hyperbolic and a 1D parabolic unstable PDE, and evaluate its performance in terms of overshoot, convergence rate, and steady-state error, compared to the backstepping controller, standard SAC, and SAC embedded with a non-pretrained DeepONet.
-Simulations on both systems demonstrate that the proposed method outperforms the other three controllers.
-The robustness against model mismatch is also evaluated by changing the system coefficient to a value different from that used for control design.
-The results show that the proposed method achieves superior performance compared to the other approaches, owing to the pre-trained DeepONet’s ability to adapt to varying coefficient functions and generate appropriate control signals in response to these changes.
-Future research will explore integrating safety control with this learning-based approach.
+This paper proposes a novel neural-operator-embedded SAC architecture for PDE control, integrating rigorous control knowledge via a DeepONet approximating the backstepping controller.
+Applied to 1D hyperbolic and parabolic unstable PDEs, it is evaluated against backstepping control, standard SAC, and SAC with non-pretrained DeepONet in terms of overshoot, convergence rate, and steady-state error.
+Simulations show superior performance.
+Robustness to model mismatch (via modified system coefficients) is validated, with pre-trained DeepONet enabling adaptation to coefficient variations and generating appropriate control signals, ensuring better performance.
+Future work will explore integrating safety control with this learning-based approach.